Find the partial differential equation whose general solution is given by the arbitrary

function F(2x + 3y + 4z , x^2 + y^2 − 2) = 0


1
Expert's answer
2021-05-18T17:28:56-0400

Given function is-


"F(2x+3y+4z,x^2+y^2-2)=0"


Here, "u=2x+3y+4z,v=x^2+y^2-2"


Now differentiate f w.r.t x-


"\dfrac{df}{du}.\dfrac{du}{dx}+\dfrac{df}{dv}.\dfrac{dv}{dx}=0~~~~~~~~~-(1)"


Also differentiate w.r.t y-


"\dfrac{df}{du}.\dfrac{du}{dy}+\dfrac{df}{dv}.\dfrac{dv}{dy}=0~~~~~~~~~-(2)"


Now eleminating function f between eqn.(1) and (2) and we get the required PDE as-


"\begin{vmatrix}\dfrac{du}{dx}&\dfrac{dv}{dx}\\\\\dfrac{du}{dy}&\dfrac{dv}{dy}\end{vmatrix}=0"


"\begin{vmatrix} \dfrac{du}{dx}& 2x\\\\\dfrac{du}{dy}&2y\end{vmatrix}=0"


"2y\dfrac{du}{dx}-2x\dfrac{du}{dy}=0"


Hence The required PDE is "2y\dfrac{du}{dx}-2x\dfrac{du}{dy}=0"